Calculate the heat required to vaporize 7.87 g of benzene at its normal boiling point.

Heat of vaporization (benzene) = 30.7 kJ/mol
1
Expert's answer
2017-12-10T06:53:07-0500
Find the amount of benzene (Mbenzene = 78 g/mol)

ν = 7.87/78 = 0.1 (mole)

Find the heat required:

Q = 0.1 × 30.7 = 3.07 (kJ)